OpenStudy (anonymous):

What is the 10th term of the geometric sequence –1, 3, –9, …?

OpenStudy (lalaly):

first term=-1 ratio=-3 10th term= -1*(-3)^(10-1) = -1*(-3)^9

OpenStudy (lalaly):

your pic is scary

OpenStudy (lalaly):

:P

OpenStudy (anonymous):

In general the \( n- \)th term of geometric sequence is given by \( T_n = a \times r^{n-1} \) where \( a \) is the first term and \( r \) is the common ratio.

OpenStudy (lalaly):

yes^

OpenStudy (anonymous):

I agree lallalallla :D

OpenStudy (anonymous):

Lol it's not scary it's cool :P but thanks with the problem

OpenStudy (anonymous):

It's not cool in any universe.

OpenStudy (lalaly):

haha... your welcome :D

OpenStudy (anonymous):

To each their own.